# Directional testing for one-way MANOVA in divergent dimensions

### v3

@@ 3.2 Heteroscedastic one-way MANOVA
We also consider the likelihood ratio test and its modifications proposed by skovgaard:2001. The likelihood ratio statistic (1) takes the form
@@
The statistic $W$ has approximate $\chi^{2}_{d}$ null distribution with degrees of freedom d=p(g1)d=p(g-1), when $p$ is fixed. The expression of the correction factor $\gamma$, given in (S1), for hypothesis (12) in skovgaard:2001’s modifications (3) becomes
@@
If $g=2$, hypothesis (12) reduces to the multivariate Behrens-Fisher problem. Then, we can compare the directional test also with the procedure proposed by nel1986, which was shown to maintain a reasonable empirical type I error. The test by nel1986 is based on the quantity
@@
See rencher1998multivariate for more details.
@@ 4.1 Homoscedastic one-way MANOVA
The performance of the directional test for hypothesis (11) in the high dimensional multivariate normal framework is here assessed via Monte Carlo simulations based on $10,000$ replications. The exact directional test is compared with the chi-square approximations for $W$, $_{BC}, WW^{*}$, $W^{**}$, and with the normal approximation for the central limit theorem test proposed by He2020B, specifically for the high dimensional setting. The six tests are evaluated in terms of empirical size.
@@
Figure 2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (11) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left and right panels correspond to $n_{i}=100,500$, respectively.
The empirical size, i.e. the actual probability of type I error, at nominal level $\alpha=0.05$ based on the null distribution of the various statistics is reported in Figure S2. The directional $p$-value performs very well across all different values of $p$, confirming the exactness result in Theorem 1, while the central limit theorem test is less accurate when $p$ is small. The test based on $W$ breaks down in all settings. Instead, the Bartlett corrected test $W_{BC}$ proves accurate for moderate values of $p$, as seen in He2020B. However, the chi-square approximations for $_{BC}, WW^{*}$ and $W^{**}$ get unreliable as $p$ grows. Table 1 reports the empirical size in some extreme settings with $n_{i}\leq p\leq\sum_{i=1}^{g}n_{i}-g-1$. The results show that the chi-square approximations for $W$, $_{BC}, WW^{*}$ and $W^{**}$ do not work in such extreme settings, while the directional test is still accurate. The central limit theorem test, instead, becomes too liberal with large $p$.
Table 1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (11) with $g=3$, $p=\kappa n_{i}$ and $n_{i}=100$, at nominal level $\alpha=0.05$.
@@ 4.2 Heteroscedastic one-way MANOVA
The performance of the directional test for hypothesis (12) in a moderate dimensional multivariate normal framework with $p=O(n_{i}^{\tau})$, $\tau\in(0,1)$ is here evaluated via Monte Carlo simulations based on 10,000 replications. The directional test is again compared with the chi-square approximations for the likelihood ratio test and its modifications proposed by skovgaard:2001. When $g=2$, we also consider the $F$-approximation for the Behrens-Fisher test $T^{*2}$ by nel1986. The different testing approaches are evaluated in terms of empirical size. The results for $g=5,30$ are reported in Section S2.3 of the Supplementary Material and are in line with the ones discussed below.
@@
Figure 3: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (12) with $g=2$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left and right panels correspond to $n_{i}=100,500$, respectively.
Figure S10 reports the empirical size at the nominal level $\alpha=0.05$ under the null hypothesis for various statistics. The directional test is always more reliable than its competitors in terms of the empirical size, even if in this framework it is not exact. skovgaard:2001’s modifications are not as accurate when $p$ is large. Moreover, the likelihood ratio test and the Behrens-Fisher test break down quickly when $p=\lceil n_{i}\rceil^{\tau}$, even for moderately large values of $\tau$. Additional simulation results are shown in Table 2 for different structures of the covariance matrices: (I) identity matrix, $\Lambda_{i}^{-1}=I_{p}$; (II) compound symmetric matrix, $\Lambda^{-1}_{i}=(1-\rho_{i})I_{p}+\rho_{i}1_{p}1_{p}^{\top}$, with the same values of $\rho_{i}$ as in the autoregressive structure. Table 2 reports the empirical size at the nominal level $\alpha=0.05$ for $n_{i}=100$ and $p=\lceil n_{i}\rceil^{\tau}$, with covariance structures (I) and (II) and $g=2,30$. As expected, the likelihood ratio test performs in general very poorly, especially for larger values of $g$. skovgaard:2001’s modifications are much more accurate but not as much as the directional test, whose excellent performance is confirmed. When $g=2$, the approximate solution for the Behrens-Fisher test seems reliable in the identity covariance matrix case, but not under compound symmetry.
Table 2: Empirical size of the directional test (DT), likelihood ratio test (LRT), two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) (only for $g=2$) for hypothesis (12) at nominal level $\alpha=0.05$, with $n_{i}=100$ and $p=n_{i}^{\tau}$.
@@
5 Applications
The dataset is taken from a study on how $p=5$ characteristics of the blood (red cell count, hemoglobin concentration, hematocrit, white cell count, and plasma ferritin concentration) vary with sport, body size, and sex of the athletes (telford1991sex). Our goal is to investigate whether the blood characteristics of athletes differ between endurance-related events and power-related events. The data can be found in the function ais of the R (Rcode) package DAAG (daag24). We performed several tests on this dataset to study the difference in blood parameters between athletes performing endurance-related and power-related sports. Typically, the conclusions drawn by the directional approach align with those from its best competitors. As an illustration, we report here some results involving two hypotheses: (i) testing the equality of blood characteristics between athletes competing in an endurance-related event (400-meter track event, $T_{400}$) and those in a power-related event (sprint, $T_{Sprint}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{T_{Sprint}}$; and (ii) testing the equality of blood characteristics for three groups of athletes that perform endurance-related sports ($Row$, SwimSwim, and $T_{400}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}$. Here we focus exclusively on data from male athletes as each blood variable is significantly different between genders (telford1991sex). We also consider the variable plasma ferritin concentration on the logarithmic scale. The sample sizes of $T_{400}$, $T_{Sprint}$, $Row$ and $Swim$ groups are respectively 18, 11, 15, 13. All the testing procedures under analysis can then be used to perform one-way MANOVA, and the corresponding $p$-values are shown in Table 3.
@@ 5 Applications
Table 3: $p$-values of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
@@
Table 4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
@@
For hypothesis (ii), all the tests used under heteroscedasticity in Table 3 give $p$-values directional: 0.097; likelihood ratio: 0.062; Skovgaard’s: 0.152 and 0.158) concluding that differences in mean blood characteristics among endurance-related events ($T_{400}$, $Row$ and $Swim$) for male athletes are not statistically significant at a 5% level. In the homoscedastic case, only the likelihood ratio test yields a significant $p$-value (0.045). Conversely, the other test statistics have $p$-values (directional: 0.092; central limit theorem: 0.069; Bartlett corrected: 0.084; Skovgaard’s: 0.101 and 0.105) that agree with the hypothesis of no significant difference among the three groups. Even in this case, a simulation study can be performed in order to verify the reliability of the considered methods. Under the null hypothesis, we generate $10,000$ samples with sizes $n_{T_{400}}=18$, $n_{Row}=15$ and $n_{Swim}=13$ from a normal distribution with mean vector $\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}=(n_{T_{400}}\bar{y}_{T_{400}}+n_{Row}\bar{y}_{Row}+n_{Swim}\bar{y}_{Swim})/(n_{T_{400}}+n_{Row}+n_{Swim})$ where $\bar{y}_{T_{400}}$, $\bar{y}_{Row}$ and $\bar{y}_{Swim}$ are the group sample means, and covariance matrix $\Lambda^{-1}=\hat{\Lambda}_{0}^{-1}$, the pooled sample covariance matrix under the homoscedasticity, and covariance matrices $\Lambda^{-1}_{i}=\hat{\Lambda}_{i}^{-1}$, sample covariance matrices under the heteroscedasticity. Simulation results are reported in Table 4. In the homoscedastic case, the two Skovgaard’s modifications are conservative. The central limit theorem test and the likelihood test, on the other hand, have slightly inflated empirical size. Hence, the $p$-values of these four methods are not reliable in this scenario. In contrast, the directional test and the Bartlett corrected test well control the type I error. In the heteroscedastic case, the likelihood ratio test is overly inflated, whereas the remaining procedures, including the directional test, exhibit a good performance.
@@ 6 Conclusions
In this paper, we have developed the directional test for one-way MANOVA problems when the data dimension is comparable with the sample size. The directional $p$-value has been proved exactly uniformly distributed in a high dimensional setting, provided that $\sum_{i=1}^{g}n_{i}\geq p+g+1$, when testing the equality of normal mean vectors with identical covariance matrix. This finding is supported by our numerical studies. In addition, simulations in moderate dimensional scenarios in the heteroscedastic one-way MANOVA framework indicate that the directional test outperforms its competitors in terms of empirical null distribution, even in the more general setting with different covariance matrices. In general, all the considered approaches rely on the assumed multivariate normal model and are not expected to be reliable if that is misspecified. Yet, our numerical results show that these deviations from normality do not affect much the behaviour of the various tests and identify the directional test as the preferable solution even under these model misspecifications. While formal conditions for the validity of the various methods in this scenario could perhaps be developed using recent results on the accuracy of the saddlepoint approximation in moderate dimensional regimes (tang2021laplace), we believe further work is needed in this area to refine these methods under more complex settings. It is also worth mentioning that, given the inconclusive results of the power analysis of likelihood-based tests for high dimensional normal distributions included in huang2021directional, here we have chosen not to compare the performance of the tests under a few arbitrary alternative hypotheses. Further investigations have showed that all the likelihood-based solutions examined are reasonably robust to the misspecification of the assumed multivariate normal model, although different and/or more substantial deviations from normality could be considered in future research for a more complete picture.
@@
S1.1 Correction factor $\gamma$ in skovgaard:2001’s modifications
The general expression for $\gamma$ appearing in skovgaard:2001’s modified versions of W (3) in the main text can be written as
@@
where $J_{\lambda\lambda}$ and $I_{\lambda\lambda}$ are the blocks of observed and expected information matrices, respectively, relative to the nuisance parameter $\lambda$. For calculating the $p$-value, the quantity (S14) is evaluated at $s=s^{0}=0_{q}$, corresponding to the observed data point $y^{0}$. In particular, if $\lambda$ is a component or a linear function of the canonical parameter, as in Section 3.1 in the main text, the last factor in (S14) equals 1.
@@
When testing hypothesis (8) in the main text, the saddlepoint approximation (salvan:1997, Section 10) and the correction factor $\gamma$ in (3) require the computation of the observed information matrix $J_{\varphi\varphi}(\varphi)$. This has the block form
@@
and $J_{\text{vech}(\Lambda)\text{vech}(\Lambda)}=\sum_{i=1}^{g}\frac{n_{i}}{2}D_{p}^{\top}\Lambda^{-1}(I_{p}+2\xi_{i}\xi_{i}^{\top}\Lambda^{-1})\otimes\Lambda^{-1}D_{p},$ where $\otimes$ denotes the Kronecker product (see, for instance, Lauritzen:1996). Then, the determinant of $J_{\varphi\varphi}(\varphi)$ can be computed as
@@
According to the Sherman-Morrison formula (see also sartori:2019Ftest, Section 4.2), we have that
@@
This section reports additional empirical result for homoscedastic one-way MANOVA in the multivariate normal framework. We compare the performance of exact directional test (DT) with other five approximate approaches: the central limite theorem test (CLT), log-likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2). The six tests are evaluated in terms of empirical size. The simulation results are computed via Monte Carlo simulation based on 10,000 replications.
Figure S1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to $n_{i}=1000$, respectively.
S3.1 Empirical results for high dimentional setup
Groups of size $n_{i},i\in\{1,\dots,g\}$, are generated from a $p$-variate standard normal distribution $N_{p}(0_{p},I_{p})$ under the null hypothesis. For each simulation experiment, we show results for $p=\kappa n_{i}$ with $\kappa\in\{0.05,0.1,0.3,$ 0.5,0.7,0.9}0.5,0.7,0.9\} and $n_{i}=1000$. Throughout, we set $g=3$ and $n_{1}=n_{2}=n_{3}$. Figure S1 shows the empirical size of the directional test, central limit theorem test, likelihood ratio test, Bartlett corrected test and two skovgaard:2001’s modifications for hypothesis (8) with $g=3$, at nominal level $\alpha=0.05$. The directional $p$-value performs better than the other. The log-likelihood ratio test is disastrous in the large sample size setting with $n_{i}=1000$.
Table S1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $n_{i}=100,500,1000$, at nominal level $\alpha=0.05$.
@@
Figure S2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to $n_{i}=100,500,1000$, respectively ($g=3$).
S3.3 Empirical results for large number $g$ of groups
Table S2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$, $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$
@@
Figure S3: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
S3.4 Empirical results for the setup of He2020B
@@
Table S3: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}100=300$, at nominal level $\alpha=0.05$.
@@
Figure S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}100=300$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
Table S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}500=1500$, at nominal level $\alpha=0.05$.

### v4

@@ 3.2 Heteroscedastic one-way MANOVA
We also consider the likelihood ratio test and its modifications proposed by Skovgaard (2001). The likelihood ratio statistic (1) takes the form
@@
The statistic $W$ has approximate $\chi^{2}_{d}$ null distribution with degrees of freedom d=p(g1)d=p(g-1), when $p$ is fixed. The expression of the correction factor $\gamma$, given in (S1), for hypothesis (12) in Skovgaard’s modifications (3) becomes
@@
If $g=2$, hypothesis (12) reduces to the multivariate Behrens-Fisher problem. Then, we can compare the directional test also with the procedure proposed by Nel and Merwe (1986), which was shown to maintain a reasonable empirical type I error. The test by Nel and Merwe (1986) is based on the quantity
@@
See Rencher (1998, Section 3.9) for more details.
@@ 4.1 Homoscedastic one-way MANOVA
The performance of the directional test for hypothesis (11) in the high dimensional multivariate normal framework is here assessed via Monte Carlo simulations based on $10,000$ replications. The exact directional test is compared with the chi-square approximations for $W$, $_{BC}, WW^{*}$, $W^{**}$, and with the normal approximation for the central limit theorem test proposed by He et al. (2021), specifically for the high dimensional setting. The six tests are evaluated in terms of empirical size.
@@
Figure 2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (11) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left and right panels correspond to $n_{i}=100,500$, respectively.
The empirical size, i.e. the actual probability of type I error, at nominal level $\alpha=0.05$ based on the null distribution of the various statistics is reported in Figure S2. The directional $p$-value performs very well across all different values of $p$, confirming the exactness result in Theorem 1, while the central limit theorem test is less accurate when $p$ is small. The test based on $W$ breaks down in all settings. Instead, the Bartlett corrected test $W_{BC}$ proves accurate for moderate values of $p$, as seen in He et al. (2021). However, the chi-square approximations for $_{BC}, WW^{*}$ and $W^{**}$ get unreliable as $p$ grows. Table 1 reports the empirical size in some extreme settings with $n_{i}\leq p\leq\sum_{i=1}^{g}n_{i}-g-1$. The results show that the chi-square approximations for $W$, $_{BC}, WW^{*}$ and $W^{**}$ do not work in such extreme settings, while the directional test is still accurate. The central limit theorem test, instead, becomes too liberal with large $p$.
Table 1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (11) with $g=3$, $p=\kappa n_{i}$ and $n_{i}=100$, at nominal level $\alpha=0.05$.
@@ 4.2 Heteroscedastic one-way MANOVA
The performance of the directional test for hypothesis (12) in a moderate dimensional multivariate normal framework with $p=O(n_{i}^{\tau})$, $\tau\in(0,1)$ is here evaluated via Monte Carlo simulations based on 10,000 replications. The directional test is again compared with the chi-square approximations for the likelihood ratio test and its modifications proposed by Skovgaard (2001). When $g=2$, we also consider the $F$-approximation for the Behrens-Fisher test $T^{*2}$ by Nel and Merwe (1986). The different testing approaches are evaluated in terms of empirical size. The results for $g=5,30$ are reported in Section S2.3 of the Supplementary Material and are in line with the ones discussed below.
@@
Figure 3: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (Nel and Merwe, 1986), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (12) with $g=2$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left and right panels correspond to $n_{i}=100,500$, respectively.
Figure S10 reports the empirical size at the nominal level $\alpha=0.05$ under the null hypothesis for various statistics. The directional test is always more reliable than its competitors in terms of the empirical size, even if in this framework it is not exact. Skovgaard’s modifications are not as accurate when $p$ is large. Moreover, the likelihood ratio test and the Behrens-Fisher test break down quickly when $p=\lceil n_{i}\rceil^{\tau}$, even for moderately large values of $\tau$. Additional simulation results are shown in Table 2 for different structures of the covariance matrices: (I) identity matrix, $\Lambda_{i}^{-1}=I_{p}$; (II) compound symmetric matrix, $\Lambda^{-1}_{i}=(1-\rho_{i})I_{p}+\rho_{i}1_{p}1_{p}^{\top}$, with the same values of $\rho_{i}$ as in the autoregressive structure. Table 2 reports the empirical size at the nominal level $\alpha=0.05$ for $n_{i}=100$ and $p=\lceil n_{i}\rceil^{\tau}$, with covariance structures (I) and (II) and $g=2,30$. As expected, the likelihood ratio test performs in general very poorly, especially for larger values of $g$. Skovgaard’s modifications are much more accurate but not as much as the directional test, whose excellent performance is confirmed. When $g=2$, the approximate solution for the Behrens-Fisher test seems reliable in the identity covariance matrix case, but not under compound symmetry.
Table 2: Empirical size of the directional test (DT), likelihood ratio test (LRT), two Skovgaard’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) (only for $g=2$) for hypothesis (12) at nominal level $\alpha=0.05$, with $n_{i}=100$, $p=n_{i}^{\tau}$, $\Lambda_{i}^{-1}=I_{p}$ (I) and $\Lambda^{-1}_{i}=(1-\rho_{i})I_{p}+\rho_{i}1_{p}1_{p}^{\top}$(II), with the $\rho_{i}$ chosen as an equally-spaced sequence from 0.1 to 0.9 of length $g$.
@@
5 Application
We consider two real-data applications of the proposed methodology. One regards microarray data of patients suffering from breast cancer, and can be found in Section S5 of the Supplementary Material, while the other is described in detail here. The dataset is taken from a study on how $p=5$ characteristics of the blood (red cell count, hemoglobin concentration, hematocrit, white cell count, and plasma ferritin concentration) vary with sport, body size, and sex of the athletes (Telford and Cunningham, 1991). Our goal is to investigate whether the blood characteristics of athletes differ between endurance-related events and power-related events. The data can be found in the function ais of the R package DAAG (Maindonald and Braun, 2024). We performed several tests on this dataset to study the difference in blood parameters between athletes performing endurance-related and power-related sports. Typically, the conclusions drawn by the directional approach align with those from its best competitors. As an illustration, we report here some results involving two hypotheses: (i) testing the equality of blood characteristics between athletes competing in an endurance-related event (400-meter track event, $T_{400}$) and those in a power-related event (sprint, $T_{Sprint}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{T_{Sprint}}$; and (ii) testing the equality of blood characteristics for three groups of athletes that perform endurance-related sports ($Row$, SwimSwim, and $T_{400}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}$. Here we focus exclusively on data from male athletes as each blood variable is significantly different between genders (Telford and Cunningham, 1991). We also consider the variable plasma ferritin concentration on the logarithmic scale. The sample sizes of $T_{400}$, $T_{Sprint}$, $Row$ and $Swim$ groups are respectively 18, 11, 15, 13. All the testing procedures under analysis can then be used to perform one-way MANOVA, and the corresponding $p$-values are shown in Table 3.
@@ 5 Application
Table 3: $p$-values of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two Skovgaard’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
@@
Table 4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two Skovgaard’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
@@
For hypothesis (ii), all the tests used under heteroscedasticity in Table 3 give $p$-values directional: 0.097; likelihood ratio: 0.062; Skovgaard’s: 0.152 and 0.158) concluding that differences in mean blood characteristics among endurance-related events ($T_{400}$, $Row$ and $Swim$) for male athletes are not statistically significant at a 5% level. In the homoscedastic case, only the likelihood ratio test yields a significant $p$-value (0.045). Conversely, the other test statistics have $p$-values (directional: 0.092; central limit theorem: 0.069; Bartlett corrected: 0.084; Skovgaard’s: 0.101 and 0.105) that agree with the hypothesis of no significant difference among the three groups. Even in this case, a simulation study can be performed in order to verify the reliability of the considered methods in the specific setting. Under the null hypothesis, we generate $10,000$ samples with sizes $n_{T_{400}}=18$, $n_{Row}=15$ and $n_{Swim}=13$ from a normal distribution with mean vector $\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}=(n_{T_{400}}\bar{y}_{T_{400}}+n_{Row}\bar{y}_{Row}+n_{Swim}\bar{y}_{Swim})/(n_{T_{400}}+n_{Row}+n_{Swim})$ where $\bar{y}_{T_{400}}$, $\bar{y}_{Row}$ and $\bar{y}_{Swim}$ are the group sample means, and covariance matrix $\Lambda^{-1}=\hat{\Lambda}_{0}^{-1}$, the pooled sample covariance matrix under the homoscedasticity, and covariance matrices $\Lambda^{-1}_{i}=\hat{\Lambda}_{i}^{-1}$, sample covariance matrices under the heteroscedasticity. Simulation results are reported in Table 4. In the homoscedastic case, the two Skovgaard’s modifications are conservative. The central limit theorem test and the likelihood test, on the other hand, have slightly inflated empirical size. Hence, the $p$-values of these four methods are not reliable in this scenario. In contrast, the directional test and the Bartlett corrected test well control the type I error. In the heteroscedastic case, the likelihood ratio test is overly inflated, whereas the remaining procedures, including the directional test, exhibit a good performance.
@@ 6 Conclusions
In this paper, we have developed the directional test for one-way MANOVA problems when the data dimension is comparable with the sample size. The directional $p$-value has been proved exactly uniformly distributed under the null hypothesis in a high dimensional setting, provided that $\sum_{i=1}^{g}n_{i}\geq p+g+1$, when testing the equality of normal mean vectors with identical covariance matrix. This finding is supported by our numerical studies. In addition, simulations in moderate dimensional scenarios under the heteroscedastic one-way MANOVA framework indicate that the directional test outperforms its competitors in terms of empirical null distribution, even in the more general setting with different covariance matrices. While formal conditions for the validity of the various test statistics in this scenario could perhaps be developed using recent results on the accuracy of the saddlepoint approximation in moderate dimensional regimes (Tang and Reid, 2025), we believe further work is needed in this area to refine such methods under more complex settings.
It is also worth mentioning that, given the inconclusive outcome of the power analysis of likeli-hood-based tests for high dimensional normal distributions included in Huang et al. (2022, Sect. 5.3), here we have deliberately chosen not to compare the performance of the tests under a few arbitrary alternative hypotheses. Finally, although additional numerical investigations in Sections S3.1–S3.4 and S4.3 of the Supplementary Material show that all the testing solutions examined are reasonably robust to misspecification of the underlying multivariate normal model, different and/or more substantial deviations from normality could be considered in future research for gaining a more comprehensive picture on this aspect.
@@
S1.1 Correction factor $\gamma$ in Skovgaard’s modifications
The general expression for $\gamma$ appearing in Skovgaard’s modified versions of W (3) in the main text can be written as
@@
where $J_{\lambda\lambda}$ and $I_{\lambda\lambda}$ are the blocks of observed and expected information matrices, respectively, relative to the nuisance parameter $\lambda$. For calculating the $p$-value, the quantity (S14) is evaluated at $s=s^{0}=0_{q}$, corresponding to the observed data point $y^{0}$. In particular, if $\lambda$ is a component or a linear function of the canonical parameter, as in Section 3.1 of the main text, the last factor in (S14) equals 1.
@@
When testing hypothesis (8) in the main text, the saddlepoint approximation (Pace and Salvan, 1997, Section 10) and the correction factor $\gamma$ in (3) require the computation of the observed information matrix $J_{\varphi\varphi}(\varphi)$. This has the block form
@@
and $J_{\text{vech}(\Lambda)\text{vech}(\Lambda)}=\sum_{i=1}^{g}\frac{n_{i}}{2}D_{p}^{\top}\Lambda^{-1}(I_{p}+2\xi_{i}\xi_{i}^{\top}\Lambda^{-1})\otimes\Lambda^{-1}D_{p},$ where $\otimes$ denotes the Kronecker product (see, for instance, Lauritzen, 1996). Then, the determinant of $J_{\varphi\varphi}(\varphi)$ can be computed as
@@
According to the Sherman-Morrison formula (see also McCormack et al., 2019, Section 4.2), we have that
@@
This section reports additional empirical result for homoscedastic one-way MANOVA in the multivariate normal framework. We compare the performance of exact directional test (DT) with other five approximate approaches: the central limite theorem test (CLT), log-likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2). The six tests are evaluated in terms of empirical size. The simulation results are computed via Monte Carlo simulation based on 10,000 replications.
Figure S1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to $n_{i}=1000$, respectively.
S3.1 Empirical results for the high dimentional setup
Groups of size $n_{i},i\in\{1,\dots,g\}$, are generated from a $p$-variate standard normal distribution $N_{p}(0_{p},I_{p})$ under the null hypothesis. For each simulation experiment, we show results for $p=\kappa n_{i}$ with $\kappa\in\{0.05,0.1,0.3,$ 0.5,0.7,0.9}0.5,0.7,0.9\} and $n_{i}=1000$. Throughout, we set $g=3$ and $n_{1}=n_{2}=n_{3}$. Figure S1 shows the empirical size of the directional test, central limit theorem test, likelihood ratio test, Bartlett corrected test and two Skovgaard’s modifications for hypothesis (8) with $g=3$, at nominal level $\alpha=0.05$. The directional $p$-value performs better than the other. The log-likelihood ratio test is disastrous in the large sample size setting with $n_{i}=1000$.
Table S1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $n_{i}\in\{100,500,1000\}$, at nominal level $\alpha=0.05$.
@@
Figure S2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to $n_{i}=100,500,1000$, respectively ($g=3$).
S3.3 Empirical results for a large number $g$ of groups
Table S2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$, $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$
@@
Figure S3: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
S3.4 Empirical results for the setup of He et al. (2021, Section A.3)
@@
Table S3: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}100=300$, at nominal level $\alpha=0.05$.
@@
Figure S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}100=300$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
Table S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}500=1500$, at nominal level $\alpha=0.05$.